OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

@ganeshie8 @satellite73

OpenStudy (anonymous):

@texaschic101 @mathstudent55

OpenStudy (anonymous):

@Luigi0210

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

honestly idk how to even do this im so lost and my teachers on vacation and cant help me ? whats the awnser do you think

OpenStudy (texaschic101):

x - 2y = 14 x = 2y + 14 now sub 2y + 14 in for x in the other equation x + 3y = 9 2y + 14 + 3y = 9 -- subtract 14 from both sides 2y + 3y = 9 - 14 5y = -5 y = -1 now sub -1 in for y in either equation to find x x - 2y = 14 x - 2(-1) = 14 x + 2 = 14 x = 14 - 2 x = 12 solution is (12,-1)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no none at all thank you do you mind helping me with some more?

OpenStudy (luigi0210):

Would been a bit simpler to multiply it by a negative.. x would of canceled :P

OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

thanks for helping too @Luigi0210

OpenStudy (texaschic101):

I will HELP.....we can go through it together yes..Luigi....you could also solve it that way

OpenStudy (texaschic101):

can you isolate x in your problem Country Cuban ?

OpenStudy (anonymous):

okay good i need someone to guide me through the steps show me how to slove the awnser and give me the awnser can you do that?

OpenStudy (texaschic101):

oops...isolate y in either of the equations

OpenStudy (texaschic101):

can you do that ?

OpenStudy (anonymous):

wouldnt you at a positive 2 to the negative 2? and i think i can im not sure..

OpenStudy (texaschic101):

ok...your trying to do elimination. We can do it that way, or we can use substitution. You need to decide which way you want to do this. Elimination or substitution

OpenStudy (anonymous):

elimnation

OpenStudy (texaschic101):

I can show you either way, or I can show you both ways

OpenStudy (anonymous):

elimination would be easier:) show me how u get the awnser with elimination

OpenStudy (texaschic101):

-2x + y = 1 -4x + y = -1 we want to eliminate a variable, it does not matter which one. It would be easier to eliminate y by multiplying one of the equations (it does not matter which one) by a negative 1. can you do that ?

OpenStudy (anonymous):

then i have about 4 more problems that i need help on too after this :) could you help with those also?

OpenStudy (texaschic101):

I can

OpenStudy (texaschic101):

multiply one of those equations by a negative 1

OpenStudy (texaschic101):

do you want me to just do it and you can use my work for an example ?

OpenStudy (anonymous):

so -4 times -1 = 4?

OpenStudy (texaschic101):

yes....-4 * -1 = 4

OpenStudy (anonymous):

okay then what?\

OpenStudy (anonymous):

so what would the solution be to this problem?

OpenStudy (texaschic101):

multiply everything in the equation by -1. so what is y * -1 ? and what is -1 * -1 ?

OpenStudy (texaschic101):

do you want me to do the problem and you can use it as an example ?

OpenStudy (anonymous):

-1,2

OpenStudy (anonymous):

sorry idk what it is

OpenStudy (anonymous):

and sure you could do that too if youd like

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

-2x + y = 1 -4x + y = -1 --->(-1) ------------------ -2x + y = 1 4x - y = 1 (result of multiplying by -1) -----------------add 2x = 2 x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 1 + 2 y = 3 solution is (1,3)

OpenStudy (texaschic101):

do you have any questions ? about anything that I did

OpenStudy (texaschic101):

solution to that one is (1,3)

OpenStudy (anonymous):

omg thanks @texaschic101 :)

OpenStudy (anonymous):

alright here comes the next one

OpenStudy (texaschic101):

do you understand what I did ?

OpenStudy (anonymous):

duhhhhhh

OpenStudy (texaschic101):

I am ready for the next one

OpenStudy (anonymous):

and yes i understand :)

OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (texaschic101):

x + 3y = -1 --->(-2) 2x + 2y = 6 ------------ -2x - 6y = 2 (result of multiplying by -2) 2x + 2y = 6 -----------add -4y = 8 y = -2 now sub -2 in for y in either equation to find x x + 3y = -1 x + 3(-2) = -1 x - 6 = -1 x = 6 - 1 x = 5 solution is : (5,-2)

OpenStudy (anonymous):

wow that was alot better! are you ready for another?

OpenStudy (texaschic101):

what you got Texas boy ...lol

OpenStudy (anonymous):

lol i got somethin for ya texas girl

OpenStudy (anonymous):

and do you know how to change your name on openstudy to where u dont have to make a new account ?

OpenStudy (texaschic101):

no I do not......I have 2 accounts, each with different names, just in case I need it

OpenStudy (anonymous):

oh okay could u ask some people cause i wanna change my name on here but dont know how and i dont wanna lose my fans

OpenStudy (texaschic101):

I can ask around

OpenStudy (texaschic101):

or just post that question in openstudy feedback section

OpenStudy (anonymous):

let me do that haha

OpenStudy (texaschic101):

you got any more math questions ?

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (texaschic101):

2x + y = 1 2x = -y + 1 x = -1/2y + 1/2 now sub -1/2y + 1/2 in for x in the other equation 4x + 2y = -1 4(-1/2y + 1/2) + 2y = -1 -- distribute through the parenthesis -2y + 2 + 2y = -1 -- subtract 2 from both sides 0 = -3 (incorrect) NO SOLUTION

OpenStudy (anonymous):

thanks girl:) ready for another?

OpenStudy (texaschic101):

yep

OpenStudy (anonymous):

alrighty :)

OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (texaschic101):

-2x + 4y = 18 4x - 8y = -36 -- divide by 2 2x - 4y = -18 so now we have : -2x + 4y = 18 2x - 4y = -18 ---------------add 0 = 0 (correct) when the answer comes out equal, then there is infinite solutions solution is option 2

OpenStudy (anonymous):

thats my awnser ? "-2x + 4y = 18 4x - 8y = -36"

OpenStudy (texaschic101):

yes...answer is option 2

OpenStudy (anonymous):

thanks 1 more/

OpenStudy (anonymous):

Visitors to a carnival can buy an unlimited-ride pass for $50 or an entrance-only pass for $20. In one day, 282 passes were sold for a total of $10,680. How many unlimited-ride passes were sold? 114 146 168 212 @texaschic101

OpenStudy (fatimab2000):

@ganeshie8

OpenStudy (texaschic101):

u + e = 282 --- unlimited + entrance only = 27 passes sold 50u + 20e = 10,680 -- 50 per unlimited + 20 per entrance only = 10,680 now we can solve by substitution... u + e = 282 e = 282 - u now we sub 282 - u in for e in the other equation.. 50u + 20e = 10,680 50u + 20(282 - u) = 10,680 50u + 5640 - 20u = 10,680 30u = 10,680 - 5640 30u = 5040 u = 5040/30 u = 168....so there was 168 unlimited passes sold